OpenStudy (anonymous):

Find S20 for the geometric series 4 + 12 + 36 + 108 + … 5,034,043,187 5,123,498,245` 6,973,568,800 7,321,089,348

rishavraj (rishavraj):

for geometric series \[S_n = a_1 \times (\frac{ 1 - r^n }{ 1 - r } )\]

OpenStudy (anonymous):

so what does that mean

rishavraj (rishavraj):

a_1 is first term i.e 4 r is common ratio i.e 3

rishavraj (rishavraj):

so for S_20 \[S_{20} = 4 \times \frac{ 1 - 3^{20} }{ 1 - 3 } \]

OpenStudy (mathstudent55):

Find the common ratio, r, by dividing a term by its previous term. Then substitute 4 for a1, the value you find for r for r, and 20 for n in the formula @rishavraj gave you above.

OpenStudy (anonymous):

thank you!!